A body covered a distance of 5 km along a semicircular path. The ratio of distance to displacement is

Grade:11

1 Answers

Arun
6 years ago
Dear Ankit
 
Distance travelled = 2\pi*r/2 = \pi*r
But displacement = shortest distance between end point and starting point. = 2r
Hence ratio of distance to displacement = \pi*r/2*r = \pi/2
= 22/7*2 = 11/7
 
Regards
Arun (askIITians forum expert)